# RL for Reasoning by Adaptively Revealing Rationales

### v1

@@ Appendix B Discussion on Learning Parities
The problem of learning parity functions has been extensively studied in the theory of machine learning [1, 2, 9, 4, 17, 21]. Typically, the task is defined as follows [^4] We follow standard theoretical conventions and use ±$\pm 1$ values for the bits. Equivalent results hold for ${0,1}$ bits. Moreover, since Transformers embed inputs into continuous vectors, empirical results are invariant to the bit representation.: each bit in the input is sampled independently and uniformly from the Rademacher distribution $\mathrm{Rad}(1/2)$ (i.e., $-1$ or $1$ with equal probability). The target function is a parity over a subset of bits, i.e., $\prod_{i\in S}x_{i}$ for some $S\subseteq[d]$. The size of $S$, denoted $|S|$, is called the degree of the parity. When $|S|=O_{d}(1)$ is constant (with respect to the input dimension), the problem is referred to as learning a sparse parity.
@@
Due to symmetry, we often discuss learning a parity of degree $k$ without specifying the particular subset $S$. Most learning formulations assume the degree $k$ is known while the identity of $S$ must be learned, yielding a hypothesis class of size $\binom{d}{k}$. It is well established that the difficulty of learning sparse parities increases with the degree $k$. In the statistical query (SQ) model [31], it has been shown that learning a parity of degree $k$ requires $\Omega(d^{k})$ queries, which translates to $\Omega(d^{k-1})$ samples [33].
@@
Similar lower bounds have been conjectured/shown for neural networks. In particular, fully connected networks with bounded width and depth and rotationally invariant weight initializations are conjectured to require $\Omega(d^{\max(k-1,1)})$ online gradient steps to learn a degree-$k$ parity [4]. This has been proven under specific assumptions, such as for noisy population gradient descent [2, 3]. Recently, Kou et al. [33] showed that this lower bound is tight: a variant of SGD (namely, sign-SGD) can learn the task in $O(\log d)$ iterations using batches of size $\tilde{O}(d^{k-1})$. The problem becomes more subtle in the offline setting where samples can be reused. In this case, Abbe et al. [4] conjecture the optimal sample complexity to be $\Theta(d^{\max(k/2,1)})$. We note that one can often reduce the sample complexity at the cost of using larger models, where the width/depth of the model scales with $d$ [17].
@@
In our proposed chain-of-parities task, an input sequence of length $L$ contains $L{-}1$ parity targets of degree 3 and one initial parity of degree 2. Learning all of these parity targets jointly with a single model may be harder (due to interference) or easier (due to shared information, such as the influence of previous bit), depending on the model architecture—particularly positional embeddings in Transformers. A simplifying assumption is to treat the learning of each parity target as independent. Under this assumption, the problem reduces to learning parities of degree 3, for which, based on the discussions above, one would not expect learnability using only $\tilde{\Theta}(L)$ samples. In contrast, the degree-2 parity should be learnable with $\tilde{\Theta}(L)$ samples [21, 33].
@@
We now explain why AdaBack can succeed where SFT and RL fail. Consider a scenario in which the first $L{-}1$ steps of the solution, i.e., $Y_{1},Z_{1},\ldots,Y_{L-1},Z_{L-1}$, are provided, and the model must complete $Y_{L},Z_{L}$. In this setup, there are two valid completions: $(Y_{L},Z_{L})$ and its complement $(Y_{L}^{\prime},Z_{L}^{\prime})$, where $x^{\prime}$ denotes the bitwise complement of $x$. The fact that the same prefix $X_{1},\ldots,X_{L},Y_{1},\ldots,Z_{L-1}$ admits two such completions indicates that if $Z_{L}$ is a parity of the previous coordinates, then $Y_{L}$ lies in the support of this parity. This effectively reduces the problem of learning a degree-3 parity function for $Z_{L}$ to that of learning a degree-2 parity function (since one bit is already revealed)—a task that is learnable with $\tilde{\Theta}(L)$ samples. This introduces a setting where AdaBack is able to learn the task, while the combination of SFT and standard RL fails.
@@ Appendix C Method Details
To address this, we introduced a small amount of supervision-free training into the curriculum: with 10% probability, we sample $\rho_{t}^{(i)}=0$ directly, while using the uniform interval sampling with the remaining 90% probability. This stochastic exposure to 00-portion training helped close the train-test gap and accelerated convergence in the early phases of AdaBack training.
@@ Appendix D Experiment Details
We used the GRPO algorithm for all RL experiments, with 8 rollouts (generations) per input sample. Unless otherwise noted, we trained all models for at least 10,000 iterations to ensure convergence to long-term behavior. Exploring different learning rates, we found $lr=10^{-6}$ to give us stable and rather fast training. We report final test accuracy or reward as the average over the last 5 checkpoints. In rare cases where performance deteriorated at the end of training, we averaged over the last 5 iterations with increasing reward.
Our RL batch size was 256. We did not use a KL penalty or entropy regularization—neither in the loss nor in the reward—when training from base models.
@@
Appendix E Additional Figures
We show how portions and the train and test rewards evolve for GSM8k, MATH, Base-7 GSM8k, and Tensor-2 GSM8k in Figures 3, 6, 7, and 8 respectively.

### v2

@@ Appendix B Discussion on Learning Parities
The problem of learning parity functions has been extensively studied in the theory of machine learning [Abbe and Sandon, 2020, 2023, Barak et al., 2022, Abbe et al., 2023a, Edelman et al., 2024, Glasgow, 2024]. Typically, the task is defined as follows [^5] We follow standard theoretical conventions and use $\pm 1$ values for the bits. Equivalent results hold for 0,1{0,1} bits. Moreover, since Transformers embed inputs into continuous vectors, empirical results are invariant to the bit representation.: each bit in the input is sampled independently and uniformly from the Rademacher distribution $\mathrm{Rad}(1/2)$ (i.e., 1-1 or $1$ with equal probability). The target function is a parity over a subset of bits, i.e., $\prod_{i\in S}x_{i}$ for some $S\subseteq[d]$. The size of $S$, denoted $|S|$, is called the degree of the parity. When $|S|=O_{d}(1)$ is constant (with respect to the input dimension), the problem is referred to as learning a sparse parity.
@@
Due to symmetry, we often discuss learning a parity of degree $k$ without specifying the particular subset $S$. Most learning formulations assume the degree $k$ is known while the identity of $S$ must be learned, yielding a hypothesis class of size $\binom{d}{k}$. It is well established that the difficulty of learning sparse parities increases with the degree $k$. In the statistical query (SQ) model [Kearns, 1998], it has been shown that learning a parity of degree $k$ requires $\Omega(d^{k})$ queries, which translates to $\Omega(d^{k-1})$ samples [Kou et al., 2024].
@@
Similar lower bounds have been conjectured/shown for neural networks. In particular, fully connected networks with bounded width and depth and rotationally invariant weight initializations are conjectured to require $\Omega(d^{\max(k-1,1)})$ online gradient steps to learn a degree-$k$ parity [Abbe et al., 2023a]. This has been proven under specific assumptions, such as for noisy population gradient descent [Abbe and Sandon, 2023, Abbe et al., 2021]. Recently, Kou et al. [2024] showed that this lower bound is tight: a variant of SGD (namely, sign-SGD) can learn the task in $O(\log d)$ iterations using batches of size $\tilde{O}(d^{k-1})$. The problem becomes more subtle in the offline setting where samples can be reused. In this case, Abbe et al. [2023a] conjecture the optimal sample complexity to be $\Theta(d^{\max(k/2,1)})$. We note that one can often reduce the sample complexity at the cost of using larger models, where the width/depth of the model scales with $d$ [Edelman et al., 2024].
@@
In our proposed chain-of-parities task, an input sequence of length $L$ contains L1L{-}1 parity targets of degree 3 and one initial parity of degree 2. Learning all of these parity targets jointly with a single model may be harder (due to interference) or easier (due to shared information, such as the influence of previous bit), depending on the model architecture—particularly positional embeddings in Transformers. A simplifying assumption is to treat the learning of each parity target as independent. Under this assumption, the problem reduces to learning parities of degree 3, for which, based on the discussions above, one would not expect learnability using only $\tilde{\Theta}(L)$ samples. In contrast, the degree-2 parity should be learnable with $\tilde{\Theta}(L)$ samples [Glasgow, 2024, Kou et al., 2024].
@@
We now explain why AdaBack can succeed where SFT and RL fail. Consider a scenario in which the first L1L{-}1 steps of the solution, i.e., $Y_{1},Z_{1},\ldots,Y_{L-1},Z_{L-1}$, are provided, and the model must complete $Y_{L},Z_{L}$. In this setup, there are two valid completions: $(Y_{L},Z_{L})$ and its complement $(Y_{L}^{\prime},Z_{L}^{\prime})$, where $x^{\prime}$ denotes the bitwise complement of $x$. The fact that the same prefix $X_{1},\ldots,X_{L},Y_{1},\ldots,Z_{L-1}$ admits two such completions indicates that if $Z_{L}$ is a parity of the previous coordinates, then $Y_{L}$ lies in the support of this parity. This effectively reduces the problem of learning a degree-3 parity function for $Z_{L}$ to that of learning a degree-2 parity function (since one bit is already revealed)—a task that is learnable with $\tilde{\Theta}(L)$ samples. This introduces a setting where AdaBack is able to learn the task, while the combination of SFT and standard RL fails.
@@ Appendix C Method Details
To address this, we introduced a small amount of supervision-free training into the curriculum: with 10% probability, we sample $\rho_{t}^{(i)}=0$ directly, while using the uniform interval sampling with the remaining 90% probability. This stochastic exposure to 0-portion training helped close the train-test gap and accelerated convergence in the early phases of AdaBack training.
@@
Choice of the reward threshold $\tau$
Curriculum learning often requires careful hyperparameter tuning—such as deciding how long to train at each stage and with what optimizer parameters or learning rate. AdaBack simplifies this process by collapsing these decisions into a single parameter: the reward threshold $\tau$. Intuitively, $\tau$ controls the average reward the model receives; the binary search mechanism ensures that the supervision is adapted to maintain this target. We experimented with several values of $\tau$ and found the final results to be largely robust, except near extreme values (i.e., $\tau\approx 0$ or $\tau\approx 1$). We chose $\tau=0.5$ in our main experiments, as this maximizes reward variance, following the insights of Razin et al. [2024] on avoiding vanishing gradients for PPO updates.
Sensitivity to the Reward Threshold
We evaluate AdaBack with reward threshold values $\tau\in\{0.1,0.2,\dots,0.9\}$ for RL training of the finetuned Llama3.2-1B on GSM8k. Figures 5 and 6 report training reward, test accuracy, and the evolution of the average revealed portion over the course of approximately 1212k RL iterations.
It can be observed that the learning dynamics is generally similar for all the considered thresholds. It is only for high thresholds such as $\tau=0.8$ and $\tau=0.9$ that a different behavior is observed at the beginning of training. These higher thresholds result in a higher supervision portion and train reward and slightly worse test reward at the beginning of training. [^6] Note that these thresholds reveal a larger part of the answer at the beginning, hence, it is natural that the test reward is comparatively lower at the beginning. Nevertheless, as training continues, it can be observed that the final performance, train/test rewards, and portions converge to similar values, becoming essentially indistinguishable. This indicates that AdaBack is highly robust to the choice of $\tau$. Given this general robustness, we take the value $\tau=0.5$ in our main experiments which aims to maximize the reward variance, consistent with the theoretical studies of Razin et al. [2024].
Figure 5: Training reward (left) and test accuracy (right) across different AdaBack reward thresholds. Although there is some difference at the begining of training for high thresholds ($\tau=0.8$ and $\tau=0.9$), the learning curves and performance become indistinguishable as training goes on.
Figure 6: Dynamics of the average revealed portion for different reward thresholds. High thresholds such as $\tau=0.8$ and $\tau=0.9$ result in a larger average revealed portion in the initial part of training. Nonetheless, as training continues, all thresholds converge to stable supervision levels. Combined with Figure 5, this suggests AdaBack’s curriculum adapts similarly across a wide range of thresholds.
@@ Appendix D Experiment Details
We used the GRPO algorithm for all RL experiments, with 8 rollouts (generations) per input sample. Unless otherwise noted, we trained all models for at least 10,000 iterations to ensure convergence to long-term behavior. To further ensure the fairness of comparisons, we continued each standard GRPO training at least as long as the corresponding AdaBack training. Exploring different learning rates, we found $lr=10^{-6}$ to give us stable and rather fast training. We report final test accuracy or reward as the average over the last 5 checkpoints. In rare cases where performance deteriorated at the end of training, we averaged over the last 5 iterations with increasing reward.
Our RL batch size was 256. We did not use a KL penalty or entropy regularization, neither in the loss nor in the reward.
For supervised-fine tuning of the base models, we used a validation set to adjust the hyperparameters and also the number of fine tuning iterations.
@@
Appendix E LLM Usage
We have used ChatGPT and Gemini models to improve text quality at phrase and paragraph level.
Appendix F Additional Figures
We show how portions and the train and test rewards evolve for GSM8k, MATH, Base-7 GSM8k, and Tensor-2 GSM8k in Figures 3, 7, 8, and 9 respectively.
@@ Appendix F Additional Figures
Figure 11: Learning dynamics of Qwen2.5-1.5B models on GSM8k. Training reward and test accuracy for both the base and instruct variants. Both models rapidly saturate the test reward after only a few iterations, exhibiting limited learning signal, mirroring the saturation effect previously reported for Llama-3.2-3B-Instruct (see Figure 10). This suggests that the dataset is effectively solved by these models during pretraining and provides limited learning signal for RL fine-tuning.